OpenStudy (anonymous):

Calc question: Find the force against the vertical end of a tank containing water weighing 62.4 lb/ft^3 as an integral with respect to y The tank end is shaped like an isosceles triangle. The base is 6ft long, and it is 8ft tall. The water only comes up to the 5ft mark

OpenStudy (anonymous):



OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

We were given the equation F=whA where F is the fluid force, w is the weight density of the water, h is the depth, and A is the area

OpenStudy (phi):

I don't understand what the shape of the tank is. I see a triangle, but tanks are 3-d ?

OpenStudy (anonymous):

The tanks are 3D, but they only want us to find the fluid force against one side of the tank

OpenStudy (phi):

if we use x for width and y for height, then something like \[ w \int \int y \ dx \ dy \]

OpenStudy (anonymous):

Why the second integral?

OpenStudy (phi):

I'm thinking a tiny area is dx * dy we might be able to get by with one integral, but we would have to think about it.

OpenStudy (anonymous):

I tried the problem and got \[\int\limits_{0}^{5}62.4(8-y)(2)(-3/8 y+3)dy\] but apparently the depth is supposed to be 5-y instead of 8-y, and I didn't understand why that is

OpenStudy (phi):

 are you doing something like this ?